OpenStudy (anonymous):

Part 2 – Estimating Solutions – Find the Estimated Solution 1. 5/7 + 8/9 = 2. 4/5 – 1/2 = 3. 2 5/8 + 2/9 = 4. 9 5/6 – 3 1/3 = 5. 17/19 + 4 =

OpenStudy (anonymous):

1. (5 / 7) + (8 / 9) = 1.6031746 3. 2 5/8 + (2 / 9) = 2.84722222 5. (17 / 19) + 4 = 4.89473684

OpenStudy (anonymous):

2. 0.8 - 0.5 = Answer >> 3 4. 9.83333333 - 3.33333333 = Answer >> 6.5 i calculate it for number 4. like 9.8 - 3.3 = 6.5 because 9 5/6 = 9.8 and 3 1/3 = 3.3 and i subtract it.

OpenStudy (anonymous):

1. 1.6 3. 2.8 5. 4.8 make it short

OpenStudy (anonymous):

show work for all

OpenStudy (anonymous):

???

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

show work!